• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 01:32:30 on May16,2008

    Entry number 238154

    This entry is a followup to: 238131

    keyword=Current monitor coeffs updated

    Coefficients were incorrect since hacsbc2 was rebooted. I changed them back to what was used during pi0. Looks pretty close.

    Just FYI, I think the currents in the start-of-run halog entries were also off by a factor of two. I don't know if updating the coeffs in the GUI will fix that.

    It looks like the currents logged in the end-of-run entries are computed correctly.


    FIGURE 1